OpenStudy (anonymous):

Use the following to find sin2x, cos2x, and tan2x Given: cos x = 7/25, for 3pi/2

OpenStudy (jdoe0001):

\(\bf cos(x)=\cfrac{7}{25}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{a=7}{c=25}\\ \quad \\ \textit{using the pythagorean theorem }c^2=a^2+b^2\implies \pm \sqrt{c^2-a^2}={\color{blue}{ b}}\) once you find "b", then you'd have all sides of the angle use that to get the identities

OpenStudy (jdoe0001):



OpenStudy (anonymous):

x lies in 4 th quadrant. sinx and tan x are negative.  sin 2x=2sin x cos x=? \[\cos2x=\sqrt{1-\sin ^{2}(2x)}\] \[\tan 2x=\frac{ \sin 2x }{ \cos 2x }=?\] or you can find tan 2x this way. \[\tan 2x=\frac{ 2\tan x }{ 1-\tan ^{2}x }\] =?

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

yw